if x+y+1=3xy,where x (not =)to y,determine value of x+y+1?

Grade:12th pass

1 Answers

Harsh Patodia IIT Roorkee
7 years ago
There can be lot of possibilties
x=0
then y= -1

x= 1
y=1 (Not to be taken as xnot equal to y)

x=2
y+1= 6y so y=1/5

Like this other cases can be taken

So x+y+1 can have multiple values

I think some info is missing like x and y should be integers.